# Maintenance Optimization for Asset Networks with Unknown Degradation Parameters

### v2

@@ 5.1 Performance analysis of heuristic solution approaches
For the instances of the simulation study, the two-threshold control limit is optimized using a two-step simulation-based process. Given that all component replacements draw from the same pool of components, we can exploit this symmetry by assuming identical PM and OPM thresholds for each asset, i.e., $\tau_{m}^{\textrm{PM}}\equiv\tau^{\textrm{PM}}$ and $\tau_{m}^{\textrm{OPM}}\equiv\tau^{\textrm{OPM}}$. Initially, the PM threshold is optimized without OPM. Once this optimal PM threshold $\tau_{*}^{\textrm{PM}}$ is established, we focus on determining the optimal OPM threshold $\tau_{*}^{\textrm{OPM}}$ under the PM threshold $\tau_{*}^{\textrm{PM}}$. This sequential approach not only simplifies the optimization process but also significantly reduces computational complexity. The optimization procedure for both instances I.1 and I.2 is visualized in Figure 2, and the performance results for the two-threshold control limit $\pi^{\textbf{L}_{1}}_{\mathcal{N}}$, the reactive heuristic $\pi^{\textbf{L}_{1}}_{\mathcal{R}}$ and the integrated Bayes heuristic $\pi^{\textbf{L}_{1}}_{\mathcal{I}}$ are summarized in Table 2. As anticipated, increasing corrective maintenance costs combined with greater variability in degradation process parameters result in a more cautious PM strategy. Notably, optimizing the OPM threshold further reduces costs by $3.78\%$ for instance I.1 and $1.80\%$ for instance I.2. The two-threshold control limit outperforms the integrated Bayes heuristic, which performs inefficiently in the multi-asset scenario due to its disregard for the economic dependence.
(a) The best-found PM threshold for instance I.1 is $15$, and the corresponding optimal OPM threshold is $9$.
(b) The best-found PM threshold for instance I.2 is $13$, and the corresponding optimal OPM threshold is $9$.
@@ 5.2 Improving heuristic solutions through approximate policy iteration
Gen 0
$\pi_{0}^{\textbf{L}}$ $\pi^{\textbf{L}_{1}}_{\mathcal{N}}$ $\pi^{\textbf{L}_{1}}_{\mathcal{R}}$ $\pi^{\textbf{L}_{1}}_{\mathcal{N}}$ $\pi^{\textbf{L}_{1}}_{\mathcal{R}}$
@@
Gen 1
$J\big{(}\pi^{\textbf{L}_{2}}_{\theta_{1}}(f^{\textbf{L}_{2}}_{1}(h))\big{)}$ $21.585\pm 0.007$ $34.225\pm 0.010$ $18.487\pm 0.010$ $28.208\pm 0.016$
@@
Gen 2
$J\big{(}\pi^{\textbf{L}_{2}}_{\theta_{2}}(f^{\textbf{L}_{2}}_{1}(h))\big{)}$ $21.539\pm 0.006$ $25.138\pm 0.008$ $18.148\pm 0.010$ $20.523\pm 0.012$
@@
Gen 3
$J\big{(}\pi^{\textbf{L}_{2}}_{\theta_{3}}(f^{\textbf{L}_{2}}_{1}(h))\big{)}$ $\mathbf{21.518\pm 0.006}$ $24.258\pm 0.007$ $\mathbf{18.097\pm 0.010}$ $19.674\pm 0.011$
@@
Gen 0
$\pi_{0}^{\textbf{L}}$ $\pi^{\textbf{L}_{1}}_{\mathcal{N}}$ $\pi^{\textbf{L}_{1}}_{\mathcal{R}}$ $\pi^{\textbf{L}_{1}}_{\mathcal{N}}$ $\pi^{\textbf{L}_{1}}_{\mathcal{R}}$
@@
In conclusion, the results in Tables 3 and 4 demonstrate that initializing DCL with the two-threshold control limit $\pi^{\textbf{L}_{1}}_{\mathcal{N}}$ significantly reduces the number of iterations needed to achieve near-optimal performance in both the information settings $\textbf{L}_{1}$ and $\textbf{L}_{2}$. The best-found neural network policies using the open-loop feedback approach $)\pi^{\textbf{L}_{2}}_{\theta}(f^{\textbf{L}_{1}}_{2}(\tilde{h}$)) achieves similar performance as the best-found neural network policies $)\pi^{\textbf{L}_{1}}_{\theta}(f^{\textbf{L}_{1}}_{3}(\tilde{h}$)) that are directly trained on the BMDP. The improvement of the best-found policy in setting $\textbf{L}_{1}$ over the optimized two-threshold control limit is $4.14\%$ (I.1) and $7.12\%$ (I.2). This suggests that the value of integration is significant and increases with the volatility of component heterogeneity and the cost ratio $\nicefrac{{C_{m}^{\textrm{CM}}}}{{C_{m}^{\textrm{PM}}}}$. However, the performance difference between the best-found policy in settings $\textbf{L}_{1}$ and $\textbf{L}_{2}$ is only $0.88\%$ (I.1) and $1.32\%$ (I.2), suggesting that the value of information is relatively small when the parameter uncertainty is managed effectively. Finally, all trained policies outperform the integrated Bayes heuristic $\pi^{\textbf{L}_{1}}_{\mathcal{I}}$, which is the current state-of-the-art, with the best-found policy improving on it by $12.17\%$ (I.1) and $10.03\%$ (I.2).
@@ 6 Case study on degradation data of interventional X-ray system filaments
In this section, we assess the performance of the approaches outlined in Section 4 on real-world degradation data of a component crucial to the functioning of medical imaging systems. Medical imaging devices, such as IXR systems, cost about one million USD, with annual maintenance expenses around 10% of the initial cost (ECRI, 2013). Over a typical 10-year lifespan, maintenance accounts for nearly half of the total ownership cost. X-ray tubes, the most expensive part of IXR systems, are critical for image-guided procedures but prone to failure, primarily due to filament wear. Filament degradation occurs as the tungsten material evaporates during use, forming a hot spot that eventually leads to failure (Covington, 1973). See Figure 4 for a schematic representation of an IXR system.
@@
To prevent unnecessary unavailability of such scanning equipment, Philips has developed a health indicator to monitor filament wear, collecting real-time degradation data from each IXR system. This generates a large data set containing real-time degradation data, recorded as time-series of filament wear, from the start of each filament’s lifespan until either failure or the present, for all IXR systems in operation. The used data set is obtained from Drent et al. (2023, Section 7) and originates from Philips, which is a manufacturer of the IXR system. In addition to manufacturing, Philips offers maintenance and service contracts to hospitals using their IXR systems. Dutch hospitals typically operate several IXR scanners across various departments.
@@ 6.1 Degradation data of X-ray tubes in IXR systems
The data set on X-ray tube degradation consists of $52$ time series representing the degradation levels of different X-ray tubes. Let $\mathcal{I}$ be the set of all X-ray tubes for which data is available, where $|\mathcal{I}|=52$. The time series for each individual X-ray tube $i\in\mathcal{I}$ is denoted as $J_{i}$. Each data point $j\in J_{i}$ in the time series is represented as a tuple $(t_{j},x_{j})_{i}$, where $t_{j}$ is the age of the X-ray tube, and $x_{j}$ is the degradation level at that age. Each tuple $(t_{j},x_{j})_{i}$ is generated when an IXR system is operated, and the time series contain between 20,000 and 300,000 data points, covering a time span of two to five years. These time series can be transformed into a series of BMDP trajectories, enabling us to directly evaluate the performance of our methods on the data.
For confidentiality reasons, the data was left-truncated and normalized. All time series start at $x_{0}=0$ for $t_{0}=0$ and end at $x_{|J_{i}|}=50$ (i.e., $\xi_{m}\equiv 50$ for all $i\in\mathcal{I}$). For each time series, the interarrival times $(t_{j}-t_{j-1})$ between consecutive data points and the degradation increments $(x_{j}-x_{j-1})$ were computed. To account for non-operational periods such as weekends, nights, or other extended downtimes, outliers in the interarrival times were removed. This allowed for the transformation of the original time series into those based on the operational age of each X-ray tube. Furthermore, several data points in the data where the image-guided procedure was deemed too short to cause significant wear on the X-ray tube were removed. Lastly, the time was normalized so that one unit of time corresponds roughly to the minimum operational time required for practical maintenance tasks, such as dispatching a service engineer to a hospital.
A statistical analysis performed by Drent et al. (2023) on the resulting data set found no rejection of the assumption that shocks follow a Poisson process, determined that damage sizes are best modeled by a geometric distribution, and revealed heterogeneity in the data set, meaning that distribution parameters for interarrival times and shock sizes vary across components.
@@ 6.2 Numerical experiments and model calibration
In this section, we outline the model calibration process, that is, estimating the hyperparameters of the degradation model. The data set introduced in the previous section is representative for the typical conditions encountered by the asset manager in practice, i.e, the baseline information level $\textbf{L}_{0}$. To address the heterogeneity of the component population, we estimate the hyperparameters of the distributions $\Lambda_{m}$ and $\Phi_{m}$ for each $m\in\mathcal{M}$ from the available X-ray tube degradation data using the maximum likelihood estimation procedure provided by Drent et al. (2023, Online Appendix C). As in Section 5, we assume that all component replacements stem from the same pool of components, i.e., $\Lambda_{m}\overset{\textrm{d}}{\equiv}\Lambda$ and $\Phi_{m}\overset{\textrm{d}}{\equiv}\Phi$. We divided the data set $\mathcal{I}$ into a training set $\mathcal{I_{\text{train}}}$ with $|\mathcal{I_{\text{train}}}|=10$ and a test set $\mathcal{I_{\text{test}}}$ with $|\mathcal{I_{\text{test}}}|=42$. The training set $\mathcal{I_{\text{train}}}$ is a randomly selected subset of $\mathcal{I}$ used solely for parameter estimation, while the test set $\mathcal{I_{\text{test}}}$ is reserved for evaluating our methods. For a detailed overview of the estimated parameters and cost settings used in this case study, refer to Table 5. The instance CS.1 features a single asset without the consideration of economic dependence, which simplifies the problem to the maintenance problem introduced by Drent et al. (2023). In contrast, instances CS.2 and CS.3, which involve two and five machines respectively, more accurately reflect the complexity of a real-world hospital setting.
Instance $M$ $\xi_{m}$ $\mu_{\Lambda}$ $\textrm{CV}_{\Lambda}$ $\mu_{\Phi}$ $\textrm{CV}_{\Phi}$ $\alpha$ $\nicefrac{{1}}{{\beta}}$ $a$ $b$ $c_{m}^{\textrm{PM}}$ $c_{m}^{\textrm{CM}}$ $c^{\textrm{ST}}$ $\gamma$
CS.1 11 $50$ $1.414$ $0.157$ $0.487$ $0.234$ $40.696$ $28.779$ $8.924$ $9.405$ 1 5 0 0.99
CS.2 22 $50$ $1.414$ $0.157$ $0.487$ $0.234$ $40.696$ $28.779$ $8.924$ $9.405$ 1 5 1 0.99
CS.3 55 $50$ $1.414$ $0.157$ $0.487$ $0.234$ $40.696$ $28.779$ $8.924$ $9.405$ 1 5 1 0.99
@@
For the calibration of the heuristic solution approaches, we optimize both the two-threshold control limit heuristic and the integrated Bayes heuristic for the fitted degradation model. Note that the integrated Bayes approach is specifically optimized for the cost ratio of CS.1, where $\nicefrac{{c_{m}^{\text{CM}}}}{{c_{m}^{\text{PM}}}}=5$, and the retrieved solution is used for the instances CS.2 and CS.3 as well. See Table 6 for a detailed summary of the heuristic solution calibration results. All performance results in this section are retrieved using $10^{6}$ repetitions of length $10^{3}$ time units. The reported half-widths again represent asymptotic $95\%$ confidence intervals.
@@
Gen 0
$\pi_{0}^{\textbf{L}}$ $\pi^{\textbf{L}_{1}}_{\mathcal{N}}$ $\pi^{\textbf{L}_{1}}_{\mathcal{N}}$ $\pi^{\textbf{L}_{1}}_{\mathcal{N}}$
@@
Gen 1
$J\big{(}\pi^{\textbf{L}_{2}}_{\theta_{1}}(f^{\textbf{L}_{2}}_{1}(h))\big{)}$ $2.932\pm 0.002$ $10.564\pm 0.004$ $24.020\pm 0.006$
@@
Gen 2
$J\big{(}\pi^{\textbf{L}_{2}}_{\theta_{2}}(f^{\textbf{L}_{2}}_{1}(h))\big{)}$ $2.90\pm 0.002$ $10.471\pm 0.004$ $23.660\pm 0.006$
@@
Gen 3
$J\big{(}\pi^{\textbf{L}_{2}}_{\theta_{3}}(f^{\textbf{L}_{2}}_{1}(h))\big{)}$ $2.901\pm 0.002$ $10.516\pm 0.004$ $23.507\pm 0.006$
@@
Gen 0
$\pi_{0}^{\textbf{L}}$ $\pi^{\textbf{L}_{1}}_{\mathcal{N}}$ $\pi^{\textbf{L}_{1}}_{\mathcal{N}}$ $\pi^{\textbf{L}_{1}}_{\mathcal{N}}$
@@
The findings are mostly consistent with the results presented in Section 4. Notable is that the performance of the neural network policies $)\pi^{\textbf{L}_{2}}_{\theta}(f^{\textbf{L}_{1}}_{2}(\tilde{h}$)) using the open-loop feedback approach is significantly worse compared to that of the neural network policies $)\pi^{\textbf{L}_{1}}_{\theta}(f^{\textbf{L}_{1}}_{3}(\tilde{h}$)) directly trained on the BMDP. This decline in performance is likely due to the substantial increase in component heterogeneity, as measured by the higher CV in the distributions $\Lambda$ and $\Phi$. For instance CS.1, the best-found neural network policy in the $\textbf{L}_{1}$ setting achieves comparable performance to the integrated Bayes heuristic. Both policies are depicted in Figure 5 and exhibit a high degree of similarity.
@@
An overlap analysis of Figure 5 demonstrates a strong alignment between the two policies. In total, $13.23\%$ of the policy decisions were identical, and $65.36\%$ differed by no more than $2\%$ of the failure threshold $\xi_{m}$. These percentages increase significantly for the lower-left $20\times 20$ subgrid, which contains more frequently visited states: $47.50\%$ of values were identical, and $97.25\%$ differed by at most $2\%$ of $\xi_{m}$. These results indicate that DCL can produce policies that closely approximate the near-optimal integrated Bayes heuristic for a real-world component, especially in regions of the state space that are frequently encountered. Moreover, DCL produces policies with state-of-the-art performance in a multi-asset setting that more closely resembles real-world hospital conditions.

### v3

@@ 5.1 Performance analysis of heuristic solution approaches
For the instances of the simulation study, the two-threshold control limit (see Section 4.1) is optimized using a two-step simulation-based process. Given that all component replacements draw from the same pool of components, we can exploit this symmetry by assuming identical PM and OPM thresholds for each asset, i.e., $\tau_{m}^{\textrm{PM}}\equiv\tau^{\textrm{PM}}$ and $\tau_{m}^{\textrm{OPM}}\equiv\tau^{\textrm{OPM}}$. Initially, the PM threshold is optimized without OPM. Once this optimal PM threshold $\tau_{*}^{\textrm{PM}}$ is established, we focus on determining the optimal OPM threshold $\tau_{*}^{\textrm{OPM}}$ under the PM threshold $\tau_{*}^{\textrm{PM}}$. This sequential approach not only simplifies the optimization process but also significantly reduces computational complexity. The optimization procedure for both instances I.1 and I.2 is visualized in Figure 3. As anticipated, increasing corrective maintenance costs combined with greater variability in degradation process parameters results in a lower PM threshold $\tau_{*}^{\textrm{PM}}$. Optimizing the OPM threshold further reduces costs by $3.78\%$ for instance I.1 and $1.80\%$ for instance I.2. The performance results for the two-threshold control limit $\pi^{\textbf{L}_{1}}_{\mathcal{N}}$, the reactive heuristic $\pi^{\textbf{L}_{1}}_{\mathcal{R}}$ and the integrated Bayes heuristic $\pi^{\textbf{L}_{1}}_{\mathcal{I}}$ are summarized in Table 2. Notably, the two-threshold control limit outperforms the integrated Bayes heuristic, which performs poorly in the multi-asset scenario due to its disregard of economic dependence.
(a) The best-found PM and OPM thresholds for instance I.1 are $\tau_{*}^{\textrm{PM}}=15$ and $\tau_{*}^{\textrm{OPM}}=9$, respectively.
(b) The best-found PM and OPM thresholds for instance I.2 are $\tau_{*}^{\textrm{PM}}=13$ and $\tau_{*}^{\textrm{OPM}}=9$, respectively.
@@ 5.2 Improving heuristic solutions through approximate policy iteration
Gen 0 $\pi_{0}^{\textbf{L}}$ $\pi^{\textbf{L}_{1}}_{\mathcal{N}}$ $\pi^{\textbf{L}_{1}}_{\mathcal{R}}$ $\pi^{\textbf{L}_{1}}_{\mathcal{N}}$ $\pi^{\textbf{L}_{1}}_{\mathcal{R}}$
@@
Gen 1 $J\big(\pi^{\textbf{L}_{2}}_{\theta_{1}}(f^{\textbf{L}_{2}}_{1}(h))\big)$ $21.585\pm 0.007$ $34.225\pm 0.010$ $18.487\pm 0.010$ $28.208\pm 0.016$
@@
Gen 2 $J\big(\pi^{\textbf{L}_{2}}_{\theta_{2}}(f^{\textbf{L}_{2}}_{1}(h))\big)$ $21.539\pm 0.006$ $25.138\pm 0.008$ $18.148\pm 0.010$ $20.523\pm 0.012$
@@
Gen 3 $J\big(\pi^{\textbf{L}_{2}}_{\theta_{3}}(f^{\textbf{L}_{2}}_{1}(h))\big)$ $\mathbf{21.518\pm 0.006}$ $24.258\pm 0.007$ $\mathbf{18.097\pm 0.010}$ $19.674\pm 0.011$
@@
Gen 0 $\pi_{0}^{\textbf{L}}$ $\pi^{\textbf{L}_{1}}_{\mathcal{N}}$ $\pi^{\textbf{L}_{1}}_{\mathcal{R}}$ $\pi^{\textbf{L}_{1}}_{\mathcal{N}}$ $\pi^{\textbf{L}_{1}}_{\mathcal{R}}$
@@
In conclusion, the results in Tables 3 and 4 demonstrate that initializing DCL with the two-threshold control limit $\pi^{\textbf{L}_{1}}_{\mathcal{N}}$ significantly reduces the number of iterations needed to achieve near-optimal performance in both the information settings $\textbf{L}_{1}$ and $\textbf{L}_{2}$. The best-found neural network policies using the open-loop feedback approach $)\pi^{\textbf{L}_{2}}_{\theta}(f^{\textbf{L}_{1}}_{2}(\tilde{h}$)) achieve similar performance to the best-found neural network policies $)\pi^{\textbf{L}_{1}}_{\theta}(f^{\textbf{L}_{1}}_{3}(\tilde{h}$)) that are directly trained on the BMDP. The improvement of the best-found policy in setting $\textbf{L}_{1}$ over the optimized two-threshold control limit is $4.14\%$ (I.1) and $7.12\%$ (I.2). This suggests that the value of integration is significant and increases with the volatility of component heterogeneity and the cost ratio $\nicefrac{{c_{m}^{\textrm{CM}}}}{{c_{m}^{\textrm{PM}}}}$. However, the performance difference between the best-found policy in settings $\textbf{L}_{1}$ and $\textbf{L}_{2}$ is only $0.88\%$ (I.1) and $1.32\%$ (I.2), suggesting that the value of information is relatively small when the parameter uncertainty is managed effectively. Finally, all trained policies outperform the integrated Bayes heuristic $\pi^{\textbf{L}_{1}}_{\mathcal{I}}$, which is the current state-of-the-art, with the best-found policy improving on it by $12.17\%$ (I.1) and $10.03\%$ (I.2).
@@ 6 Case study on degradation data of interventional X-ray system filaments
In this section, we assess the performance of the approaches outlined in Section 4 on real-world degradation data of a component crucial to the functioning of medical imaging systems. Medical imaging devices, such as IXR systems, cost about one million USD, with annual maintenance expenses around $10\%$ of the initial cost [ECRI, 2013]. Over a typical 10-year lifespan, maintenance accounts for nearly half of the total ownership cost. X-ray tubes, the most expensive components of IXR systems, are critical for image-guided procedures but prone to failure, predominantly due to filament wear. During each imaging procedure, a high electric current is applied to the tungsten filament to generate electrons for the X-ray beam. This repeated heating causes small amounts of tungsten to evaporate, thinning the filament and forming a hotspot that eventually leads to failure [Covington, 1973]. Importantly, degradation increments occur only during X-ray exposures and only when the electric current applied to the filament is sufficiently high to cause material degradation. Consequently, degradation accumulates through a sequence of operational events associated with system usage. In our modeling framework, these usage events are interpreted as the stochastic shocks that produce incremental degradation of the filament. We refer the reader to Figure 5 for a schematic representation of an IXR system.
@@
To prevent unnecessary unavailability of such scanning equipment, Philips, which is a manufacturer of the IXR system, has developed a health indicator to monitor filament wear, collecting real-time degradation data from each IXR system. This generates a large data set containing real-time degradation data, recorded as time-series of filament wear, from the start of each filament’s lifespan until either failure or the present, for all IXR systems in operation. The data set used is obtained from Drent et al. [2023, Section 7] and originates from Philips. In addition to manufacturing, Philips offers maintenance and service contracts to hospitals using their IXR systems. Dutch hospitals typically operate several IXR scanners across various departments.
This case study serves as a realistic industrial illustration of the proposed methodology. The modeling framework and DRL-based solution approach are not specific to medical imaging systems; rather, they apply to asset networks in which degradation can be represented by stochastic increment or shock processes with learnable parameters, such as wind turbines, semiconductor equipment, or other high-value capital goods.
@@ 6.1 Degradation data of X-ray tubes in IXR systems
The X-ray tube degradation data set consists of $52$ time series, each representing the degradation level of a distinct X-ray tube. Let $\mathcal{I}$ denote the set of X-ray tubes for which data is available, i.e., $|\mathcal{I}|=52$. The time series for each X-ray tube $i\in\mathcal{I}$ is denoted as $\mathcal{J}_{i}$. Each data point in the time series $\mathcal{J}_{i}$ is represented as a tuple $(t_{j},x_{j})_{i}$, where $t_{j}$ is the age of the X-ray tube, and $x_{j}$ is the degradation level at that age. Each tuple $(t_{j},x_{j})_{i}$ is generated when an IXR system is operated, and the time series contain between $20,000$ and $300,000$ data points, covering a time span of two to five years. These time series can be transformed into a series of BMDP trajectories, enabling us to directly evaluate the performance of our methods on the data.
For confidentiality reasons, the data was left-truncated and normalized. All time series start at $x_{0}=0$ for $t_{0}=0$ and end at $x_{|\mathcal{J}_{i}|}=50$ (i.e., $\xi_{m}\equiv 50$ for all $i\in\mathcal{I}$). For each time series, the interarrival times $(t_{j}-t_{j-1})$ between consecutive data points and the degradation increments $(x_{j}-x_{j-1})$ were computed. To account for non-operational periods such as weekends, nights, or other extended downtimes, outliers in the interarrival times were removed. This allowed for the transformation of the original time series into those based on the operational age of each X-ray tube. Furthermore, several data points in the data where the image-guided procedure was deemed too short to cause significant wear on the X-ray tube were removed. Lastly, the time was normalized so that one unit of time corresponds roughly to the minimum operational time required for practical maintenance tasks, such as dispatching a service engineer to a hospital.
A statistical analysis of the resulting data set revealed that (i) there was no evidence to reject the assumption that shocks arrive according to a Poisson process, (ii) damage sizes are best modeled by a geometric distribution, and (iii) the data exhibits heterogeneity, meaning that the distribution parameters for interarrival times and shock sizes vary across components. Indeed, the estimated hyperparameters (see Table 5) indicate significant component heterogeneity. For the shock rate $\lambda_{m}$, the statistical analysis yields a mean of $1.414$ and a CV of $0.157$. Similarly, for the shock size parameter, the mean is $0.487$ with a CV of $0.234$.
@@ 6.2 Numerical experiments and model calibration
In this section, we outline the model calibration process, that is, estimating the hyperparameters of the degradation model. The data set introduced in the previous section is representative of the typical conditions encountered by the asset manager in practice, i.e., the baseline information level $\textbf{L}_{0}$. To address the heterogeneity of the component population, we estimate the hyperparameters of the distributions $\Lambda_{m}$ and $\Phi_{m}$ for each $m\in\mathcal{M}$ from the available X-ray tube degradation data using the maximum likelihood estimation procedure provided by Drent et al. [2023, Online Appendix C]. As in Section 5, we assume that all component replacements stem from the same pool of components, i.e., $\Lambda_{m}\overset{\textrm{d}}{\equiv}\Lambda$ and $\Phi_{m}\overset{\textrm{d}}{\equiv}\Phi$. This assumption reflects asset-level homogeneity and is justified because, although the data may stem from distinct machines operated in different locations, the machines are of the same type, and the operating environment for each is controlled and standardized.
We divided the data set $\mathcal{I}$ into a training set $\mathcal{I_{\text{train}}}$ with $|\mathcal{I_{\text{train}}}|=10$ and a test set $\mathcal{I_{\text{test}}}$ with $|\mathcal{I_{\text{test}}}|=42$. The training set $\mathcal{I_{\text{train}}}$ is a randomly selected subset of $\mathcal{I}$ used solely for parameter estimation, while the test set $\mathcal{I_{\text{test}}}$ is reserved for evaluating our methods. For a detailed overview of the estimated parameters and cost settings used in this case study, refer to Table 5. The instance CS.1 features a single asset without consideration of economic dependence, which simplifies the problem to the maintenance problem studied in Drent et al. [2023]. In contrast, instances CS.2 and CS.3, which involve two and five machines respectively, more accurately reflect the complexity of a real-world hospital setting.
$\mu_{\Lambda}$ $\textrm{CV}_{\Lambda}$ $\mu_{\Phi}$ $\textrm{CV}_{\Phi}$ $\alpha$ $\nicefrac{{1}}{{\beta}}$ $r$ $s$
$1.414$ $0.157$ $0.487$ $0.234$ $40.696$ $28.779$ $8.924$ $9.405$
Instance $M$ $\xi_{m}$ $c_{m}^{\textrm{PM}}$ $c_{m}^{\textrm{CM}}$ $c^{\textrm{ST}}$ $\gamma$
CS.1 11 $50$ 1 5 0 0.99
CS.2 22 $50$ 1 5 1 0.99
CS.3 55 $50$ 1 5 1 0.99
@@
For the calibration of the heuristic solution approaches, we optimize both the two-threshold control limit heuristic and the integrated Bayes heuristic for the fitted degradation model. Note that the integrated Bayes approach is specifically optimized for the cost ratio of CS.1, where $\nicefrac{{c_{m}^{\textrm{CM}}}}{{c_{m}^{\textrm{PM}}}}=5$, and the retrieved solution is used for the instances CS.2 and CS.3 as well. See Table 6 for a detailed summary of the heuristic solution calibration results. All performance results in this section are obtained from $10^{6}$ repetitions, each lasting $10^{3}$ time units. The reported half-widths again represent asymptotic $95\%$ confidence intervals.
@@
Gen 0 $\pi_{0}^{\textbf{L}}$ $\pi^{\textbf{L}_{1}}_{\mathcal{N}}$ $\pi^{\textbf{L}_{1}}_{\mathcal{N}}$ $\pi^{\textbf{L}_{1}}_{\mathcal{N}}$
@@
Gen 1 $J\big(\pi^{\textbf{L}_{2}}_{\theta_{1}}(f^{\textbf{L}_{2}}_{1}(h))\big)$ $2.932\pm 0.002$ $10.564\pm 0.004$ $24.020\pm 0.006$
@@
Gen 2 $J\big(\pi^{\textbf{L}_{2}}_{\theta_{2}}(f^{\textbf{L}_{2}}_{1}(h))\big)$ $2.90\pm 0.002$ $10.471\pm 0.004$ $23.660\pm 0.006$
@@
Gen 3 $J\big(\pi^{\textbf{L}_{2}}_{\theta_{3}}(f^{\textbf{L}_{2}}_{1}(h))\big)$ $2.901\pm 0.002$ $10.516\pm 0.004$ $23.507\pm 0.006$
@@
Gen 0 $\pi_{0}^{\textbf{L}}$ $\pi^{\textbf{L}_{1}}_{\mathcal{N}}$ $\pi^{\textbf{L}_{1}}_{\mathcal{N}}$ $\pi^{\textbf{L}_{1}}_{\mathcal{N}}$
@@
The findings are mostly consistent with the results presented in Section 4. Notable is that the performance of the neural network policies $)\pi^{\textbf{L}_{2}}_{\theta}(f^{\textbf{L}_{1}}_{2}(\tilde{h}$)) using the open-loop feedback approach is significantly worse compared to that of the neural network policies $)\pi^{\textbf{L}_{1}}_{\theta}(f^{\textbf{L}_{1}}_{3}(\tilde{h}$)) directly trained on the BMDP. This decline in performance is likely attributable to a combination of several factors: (i) the substantial increase in component heterogeneity, as reflected by the higher CV of the distributions $\Lambda$ and $\Phi$; (ii) the limited number of observations available prior to component replacement; (iii) the reduction of the posterior distribution to a single point estimate, which entails a considerable loss of information; and (iv) the risk that unrepresentative initial data biases the inferred distribution parameters, potentially leading the neural network policy to trigger premature interventions.
For instance CS.1, the best-found neural network policy in the $\textbf{L}_{1}$ setting achieves comparable performance to the integrated Bayes heuristic. Both policies are depicted in Figure 6 and exhibit a high degree of similarity.
@@
An overlap analysis of Figure 6 demonstrates a strong alignment between the two policies. Overall, $13.23\%$ of the policy decisions were identical, and $65.36\%$ differed by no more than $2\%$ of the failure threshold $\xi_{m}$. These percentages increase significantly for the lower-left $20\times 20$ subgrid, which contains more frequently visited states: $47.50\%$ of values were identical, and $97.25\%$ differed by at most $2\%$ of $\xi_{m}$. These results indicate that DCL can produce policies that closely approximate the near-optimal integrated Bayes heuristic for a real-world component, especially in regions of the state space that are frequently encountered. Moreover, DCL produces policies with state-of-the-art performance in a multi-asset setting that more closely resembles real-world hospital conditions.